If a certain number is divided by 12, the quotient, dividend and divisor added together, will amount to 64. What is the number?
Answer by JoelSchwartz(130)

The quotient is the result of division which is x/12. The dividend is the x and the the twelve is the divisor.
x/12+x+12=64
x/12+x=52
x+12x=624
x=48